List of participants at Strobl

1:

Dr. P. Bastian Universität Stuttgart, Inst. f. Computeranwendungen ICA, Pfaffenwaldring 27, D-70569 Stuttgart, Germany, bastian@ica.uni-stuttgart.de

2:

Dr. Klaus Birken Universität Stuttgart, Inst. f. Computeranwendungen ICA, Pfaffenwaldring 27, D-70569 Stuttgart, Germany, klaus@ica.uni-stuttgart.de

3:

Dr. Radim Blaheta Institute of Geonics, AS CR, Studentska 1768, 708 00 Ostrava-Poruba, Czech Republic, blaheta@ugn.cas.cz

4: Dr. Alfio Borzi  Weyer 74/2, A-5163 Mattsee, Austria, alfiob@comlab.ox.ac.uk

5:

Prof. D. Braess Ruhr-Universität Bochum, Institut f. Mathematik, Universitätsstr. 150, D-44801 Bochum, Germany, braess@num.ruhr-uni-bochum.de

6:

Dr. C. Brand Montanuniversität Leoben, Institut für Angewandte Mathematik, Franz-Josef Str. 18, A-8700 Leoben, , brand@unileoben.ac.at

7:

Prof. U. Breitschuh FH Anhalt, FB Informatik, Pfeifferstr. 25, D-39114 Magdeburg, Germany, breitsch@fmi001.mi.fh-anhalt.de

8:

Prof. C. Carstensen TH Darmstadt, FB Mathematik, Schloßgartenstr. 7, D-64289 Darmstadt, Germany, cc@mathematik.th-darmstadt.de

9: J.-B. Cazier  ETH Zürich, IIS, ETZ, Gloriastr. 35, CH-8092 Zürich, Switzerland, cazier@iis.ee.ethz.ch

10:

Dr. Michael Czajkowski Universität Düsseldorf, LS für Angewandte Mathematik, Universitätsstr. 1, D-40225 Düsseldorf, Germany, michaelc@numerik.uni-duesseldorf.de

11:

Prof. W. Dahmen RWTH Aachen, Inst. f. Geom. u. Prakt. Math., Templergraben 52, D-52056 Aachen, Germany, dahmen@igpm.rwth-aachen.de

12:

Prof. C. Douglas IBM Research Division, T.J. Watson Center, P.O.Box 218, NY 10598-0218 Yorktown Heights, USA, douglas@watson.ibm.com

13: Knut Eckstein  Universität Stuttgart, Graduiertenkolleg, Pfaffenwaldring 27, D-70569 Stuttgart, Germany,
14: Dr. J. Fuhrmann  WIAS, Mohrenstraße 39, D-10117 Berlin, Germany, fuhrmann@wias-berlin.de

15:

Stefan A. Funken Universität Hannover, Inst. für Angewandte Mathematik, Welfengarten 1, D-30167 Hannover, Germany, funken@ifam.uni-hannover.de

16: Prof. Klaus Gärtner  ETH Zürich, IIS, ETZ, Gloriastr. 35, CH-8092 Zürich, Switzerland, gaertner@iis.ee.ethz.ch

17:

Dr. Csaba Gaspar Szechenyi Istvan College, Dept. of Mathematics, P.O.Box 701, H-9077 Györ, Ungarn, csgaspar@d7.szif.hu

18:

Prof. M. Griebel Universität Bonn, Institut für Angewandte Mathematik, Wegelerstr.6, D-5315 Bonn, BRD, griebel@iam.uni-bonn.de

19:

Dr. G. Haase Johannes Kepler Universität Linz, Institut für Mathematik, Altenbergerstr. 69, A-4040 Linz, Austria, ghaase@numa.uni-linz.ac.at

20:

Prof. W. Hackbusch Christian-Albrechts-Universität, Inst.f.Informatik u.Prakt.Math., Olshausenstr. 40, D-24098 Kiel, Germany, wh@informatik.uni-kiel.de

21:

Dr. B. Heise Johannes Kepler Universität, Institut für Mathematik, Altenbergerstr. 69, A-4040 Linz, Austria, heise@numa.uni-linz.ac.at

22:

Dr. Oleg Iliev University of Erlangen-Nürnberg, LSTM, Cauerstr. 4, D-91058 Erlangen, Germany, iliev@lstm.uni-erlangen.de

23:

Dr. Volker John Otto-von-Guericke-Universität, Magdeburg, PSF 4120, D-39016 Magdeburg, BRD, Volker.John@Mathematik.Uni-Magdeburg.de

24:

Dr. M. Jung TU Chemnitz-Zwickau, Fakultät Mathematik, PSF 964, D-09107 Chemnitz, Germany, dr.michael.jung@mathemathik.tu-chemnitz.de

25:

F. Kickinger Johannes Kepler Universität Linz, Institut für Mathematik, Altenbergerstr. 69, A-4040 Linz, Austria, kick@numa.uni-linz.ac.at

26:

Dr. Bela Kiss Szechenyi Istvan College, Dept. oof Mathematics, Hedervari u. 3, H-9026 Györ, Hungaria, bkiss@d7.szif.hu

27:

Ottmar Klaas Universität Hannover, Inst.f. Baumechanik und Num.Math., Appelstr. 9a, D-30167 Hannover, BRD, ottmar@hamy.ifbnm.uni-hannover.de 28:

Dr. Petr Knobloch Otto-von-Guericke-Universität, Magdeburg, PSF 4120, D-39016 Magdeburg, BRD, knobloch@dfnserv1.urz.Uni-Magdeburg.de

29:

J. Kraus Montanuniversität Leoben, Institut für Angewandte Mathematik, Franz-Josef Str. 18, A-8700 Leoben, , kraus@unileoben.ac.at

30:

Michael Kuhn Johannes Kepler Universität Linz, Institut für Mathematik, Altenbergerstr. 69, A-4040 Linz, Austria, kuhn@numa.uni-linz.ac.at

31:

Reinhard Kulke

IMST,  Carl-Friedrich-Gauss-Str.2,  D-47475  Kamp-Lintfort,  BRD,

kulke@imst.uni-duisburg.de

32: Dr. Angela Kunoth  Weierstrass-Institut, Mohrenstr. 39, D-10117 Berlin, Germany, kunoth@wias-berlin.edu

33:

Stefan Lang Universität Stuttgart, Inst. f. Computeranwendungen ICA, Pfaffenwaldring 27, D-70569 Stuttgart, Germany, stefan@ica.uni-stuttgart.de

34:

Prof. U. Langer Johannes Kepler Universität Linz, Institut für Mathematik, Altenbergerstr. 69, A-4040 Linz, Austria, ulanger@numa.uni-linz.ac.at

35:

Dr. R. Lezius TU Chemnitz-Zwickau, Fakultät Mathematik, PSF 964, D-09107 Chemnitz, BRD, rlezius@mathemathik.tu-chemnitz.de

36:

Dr. Ewald Lindner Johannes Kepler Universität Linz, Institut für Mathematik, Altenbergerstr. 69, A-4040 Linz, Austria, lindner@numa.uni-linz.ac.at

37: Prof. S. Margenov  CLPP-BAS, Acad. G. Bontchev 25a, 1113 Sofia, Bulgaria, margenov@iscbg.acad.bg

38:

H. Matthes

TU-Chemnitz-Zwickau,  Fakultät Mathematik,  D-09107 Chemnitz,  BRD,

Holger.Matthes@Mathematik.tu-chemnitz.de

39:

M. Merkel Universität Erlangen-Nürnberg, Lehrstuhl für Technische Mechanik, Egerlandstr. 5, D-91058 Erlangen, Germany, merkel@ltm.uni-erlangen.de

40:

Prof. A. Meyer Technische Universität Chemnitz, Institut für Mathematik, Reichenhainer Str. 41, PSF 964, D-09126 Chemnitz, Germany, a.meyer@mathematik.tu-chemnitz.de

41:

Prof. Gyözö Molnarka Szechenyi Istvan College, Dept. of Mathematics, Hedervari u.3, H-9026 Györ, Hungaria, molnarka@d7.szif.hu 42: Prof. S.V.

Nepomnyaschikh

Computing Center, Lavzentiev av. 6, R-630090 Novosibirsk, Russia, svnep@comcen.nsk.su

43:

Dr. Maya Neytcheva Catholic University Nijmegen, Faculty of Math. and Informatics, Toernooiveld, NL-6525 ED Nijmegen, Nederlands, neytchev@sci.kun.nl

44:

Rainer Niekamp Universität Hannover, Inst.f. Baumechanik und Num.Math., Appelstr. 9a, D-30167 Hannover, BRD, rainer@hamy.ifbnm.uni-hannover.de

45: K.-D. Oertel  ZIB, Heibronner Str. 10, D-10711 Berlin, BRD, oertel@zib-berlin.de

46:

H. Oswald Universität Heidelberg, Inst. f. Angewandte Mathematik, INF 294, D-69120 Heidelberg, BRD, Hubertus.Oswald@IWR.uni-heidelberg.de

47:

Dr. U. Rüde

Universität Augsburg, Inst.  f.  Mathematik, D-86159 Augsburg, Germany,

ruede@informatik.tu-muenchen.de

48:

Dr. Arnold Reusken TU Eindhoven, Dept. of Math. and Comp. Science, P.O.Box 513, NL-5600 MB Eindhoven, Niederlande, wsanar@win.tue.nl

49:

Dr. Stefan Sauter Universität Kiel, Lehrstuhl Prakt. Math., D-24098 Kiel, Germany, sas@informatik.uni-kiel.de

50:

J. Schöberl Johannes Kepler Universität Linz, Institut für Mathematik, Altenbergerstr. 69, A-4040 Linz, Austria, joachim@numa.uni-linz.ac.at

51:

Prof. V. Shaidurov Computing Center of RAS, Akademgorodok, 660036 Krasnoyarsk, Russia, shidurov@cckr.krasnoyarsk.su

52:

Dr. T. Skalicky Dresden University of Technology, Institut for Fluid Mechanics, Mommsenstr. 13, D-01062 Dresden, Germany, skalicky@msmfs1.mw.tu-dresden.de

53:

Dr. Barry Smith Argonne National Laboratory, MSC Building 221, 9700 South Cass Avenue, Argonne, Illinois 60439-4844, U.S.A., bsmith@mcs.anl.gov

54:

Dr. Ivan L. Sofronov Universität Stuttgart, Math. Institut A, Pfaffenwaldring 57, D-70569 Stuttgart, Germany, sofronov@mathematik.uni-stuttgart.de

55:

Dr. Linda Stals Universität Augsburg, Institut für Mathematik, Universitätsstr.14, D-86159 Augsburg, BRD, stals@malaga.math.uni-augsburg.de 56:

Prof. G. Starke Institut für praktische Mathematik, Universität Karlsruhe, D-76128 Karlsruhe, BRD, starke@math.uni-karlsruhe.de

57:

Olaf Steinbach Universität Stuttgart, Math. Institut A, Pfaffenwaldring 57, D-70550 Stuttgart, Germany, olaf@mathematik.uni-stuttgart.de

58:

Prof. E.P. Stephan Institut für Angewandte Mathematik, Universität Hannover, Welfengarten 1, D-30167 Hannover, BRD, stephan@ifam.uni-hannover.de

59:

Dr. Rob Stevenson University Nymegen, P.O.Box 9010, NL-6500 Nymegen CL, Netherlands, stevenso@sci.kun.nl

60:

Dr. A. Tamme Johannes Kepler Universität Linz, Institut für Mathematik, Altenbergerstr. 69, A-4040 Linz, Austria, tamme@numa.uni-linz.ac.at

61:

M. Theß TU Chemnitz, Fakultät für Mathematik, PSF 964, D-09107 Chemnitz, Germany, m.thess@mathematik.tu-chemnitz.de

62:

Prof. L. Tobiska Otto-von-Guericke-Universität, Magdeburg, PSF 4120, D-39016 Magdeburg, BRD, Lutz.Tobiska@Mathematik.Uni-Magdeburg.d400.de

63:

K. Urban RWTH Aachen, Inst.für Geometrie u. Prakt.Math., Templergraben 55, D-52056 Aachen, Germany, urban@igpm.rwth-aachen.de

64:

Prof. M. Vajtersic Slovak Academy of Sciences, Institute of Technical Cybernetics, Dubravska cesta 9, 84237 Bratislava, Slovakia, kaifmava@savba.savba.sk

65:

Prof. Y. Vassilevski Russian Academy of Science, Inst. of Numerical Mathematics, Leninskij Prospect 32a, R-117334 Moscow, Russia, vasilevs@unitron1.inm.ras.ru

66:

Frank Wagner Universität Augsburg, LS für Angewandte Mathematik 1, Universitätsstr. 14, D-86159 Augburg, Germany, Frank.Wagner@Math.uni-augsburg.de

67:

Dr. Takumi Washio C& C Research Laboratories, NEC Europe Ltd., Rathausallee 10, D-53757 Sankt Augustin, BRD, washio@ccrl-nece.technopark.gmd.de

68:

Prof. G. Wittum Universität Stuttgart, Inst. f. Computeranwendungen 3, Pfaffenwaldring 27, D-70569 Stuttgart, Germany, wittum@ica.uni-stuttgart.de

69:

Ch. Wrobel Universität Stuttgart, Inst. f. Computeranwendungen 3, Pfaffenwaldring 27, D-70569 Stuttgart, Germany, christian@ica.uni-stuttgart.de 70:

Dr. Walter Zulehner Johannes Kepler Universität Linz, Institut für Mathematik, Altenbergerstr. 69, A-4040 Linz, Austria, zulehner@numa.uni-linz.ac.at